OpenStudy (anonymous):

Consider the sequence 16, –8, 4, –2, 1, ... a. Describe the pattern formed in the sequence. b. Find the next three terms

OpenStudy (anonymous):

Try to think of the different ways you could get from 16 to -8 using addition, subtraction, multiplication and division. Write them in a list, then go through the list and see which method will also get from -8 to 4.

OpenStudy (anonymous):

divide by -2 to get the next terms next 3 are (-1/2, 1/4, and -1/8)

OpenStudy (anonymous):

That was my answer as well. Great Job!

OpenStudy (anonymous):

I have another question but I think no one wants to answer it I posted it earlier...

OpenStudy (anonymous):

I'll take a look...

OpenStudy (anonymous):

Sorry, I can't help with trig. Try www.purplemath.com or www.khanacademy.org they are both great resources.

OpenStudy (anonymous):

\[\tan 4 \theta +\tan 2 \theta/ 1- \tan 4 \theta \tan 2 \theta\]